b) The odds against student X solving a Business statistics problem are 8 to 6, and odds in favour of student Y solving the problem are 14 to 16.
1. What is the chance that the problem will be solved if they both try independent of each other?
2. What is the probability that none of them is able to solve the problem?
Question 2: Marks: 1+3+2+3+1= 10
For the frequency distribution given below:
Length of Service No of employees
7.5-7.9 6
8.0-8.4 22
8.5-8.9 36
9.0-9.4 18
9.5-9.9 14
10.0-10.4 4
Calculate
(i) Mean
(ii) Standard deviation
(iii) Co-efficient of Variation
(iv) Mean Deviation (from median)
(v) Range

Question 3: Marks: 2+2=4
a) How many distinct four-digit numbers can be performed from the following integers 1, 2, 3,4,5,6 if each integer is used only once?
b) What would be the shape and name of the frequency distribution if
1. mean=median=mode
2. mean>median>mode
Question 4: Marks: 2+6=8
a) Describe the situation in which two variables are perfect positively correlated?
b) The cost of output at a factory is thought to depend on the number of units produced. Data have been collected for the number of units produced each month in the last six months, and the associated costs, as follows;
Output
(‘000s of units) X Cost
($’000) Y
2 9
3 11
1 7
4 13
3 11
5 15
Calculate the correlation coefficient and comment on your result.

Answer by stanbon(48568)

b) The odds against student X solving a Business statistics problem are 8 to 6, and odds in favour of student Y solving the problem are 14 to 16.
P(X solves) = 6/14 = 3/7
P(Y solves) = 14/30 = 7/15
---------------------------------------
1. What is the chance that the problem will be solved if they both try independent of each other? (3/7)(7/15) = 1/5
-----------------------------------------------------------
2. What is the probability that none of them is able to solve the problem?
(4/7)(8/15) = 28/105
-----------------------------------------------------------
Question 2: Marks: 1+3+2+3+1= 10
For the frequency distribution given below:
Length of Service No of employees
7.5-7.9 6
8.0-8.4 22
8.5-8.9 36
9.0-9.4 18
9.5-9.9 14
10.0-10.4 4
Calculate
(i) Mean
(ii) Standard deviation
(iii) Co-efficient of Variation
(iv) Mean Deviation (from median)
(v) Range
------------------
You should have software or a calculator to help you do this
arithmetic.
------------------
Question 3: Marks: 2+2=4
a) How many distinct four-digit numbers can be performed from the following integers 1, 2, 3,4,5,6 if each integer is used only once?
6! = 720
-----------------------------------
b) What would be the shape and name of the frequency distribution if
1. mean=median=mode
symmetric; normal
-------------------------------
2. mean>median>mode
Draw the picture and you will see it is skewed
------------------------------------------------------
Question 4: Marks: 2+6=8
a) Describe the situation in which two variables are perfect positively correlated? a line or curve
--------------------------------
b) The cost of output at a factory is thought to depend on the number of units produced. Data have been collected for the number of units produced each month in the last six months, and the associated costs, as follows;
Output
(‘000s of units) X Cost
($’000) Y
2 9
3 11
1 7
4 13
3 11
5 15
Calculate the correlation coefficient and comment on your result.
You should have software or a calculator that will do this for you.
Using a TI-83 I get this answer: r = 1
-------------
Cheers,
Stan H.